solve for x: sqaure root(3x+10) minus square root(x+3)=-1

Answer by Prithwis(166)

sqrt(3x+10) - sqrt(x+3) = -1
=> sqrt(3x+10) = sqrt(x+3) - 1
Square both sides:
(3x+10) = (x+3) + 1 - 2sqrt(x+3)
=> 2x+6 = -2sqrt(x+3)
=> x+3 = -sqrt(x+3)
Square again:
(x+3)^2 = (x+3)
=> (x+3)(x+3-1) = 0
=> (x+3)(x+2) = 0
=> x = -3 or x= -2
Answer: x = -3 or x= -2 (you will need to use both the values in the square root to check if the value within sqaure root is >= 0); In this case they are.